OpenStudy (firejay5):

How to change: x - 5

OpenStudy (lgbasallote):

\[x - 5 \le y\] ??

OpenStudy (firejay5):

yes I can't make the line under the inequality sign

OpenStudy (lgbasallote):

so what's your question?

OpenStudy (firejay5):

How would it look like in y = mx + b

OpenStudy (lgbasallote):

it can't... y = mx + b has an equal sign....you don't have an equal sign here...

OpenStudy (firejay5):

I know that

OpenStudy (lgbasallote):

so what are you asking?

OpenStudy (firejay5):

is it already in y = w/o the inequality sign

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

it's in the form \[y \ge mx + b\]

OpenStudy (lgbasallote):

but y = mx + b...no

OpenStudy (firejay5):

the next questions that I am asking will be like that, cause I need it to be y = mx + b in order to enter it into my calculator

OpenStudy (lgbasallote):

you can't turn inequalities into equalities...it's impossible

OpenStudy (firejay5):

do you understand what I mean

OpenStudy (firejay5):

Do you know what a TI 84 is????

OpenStudy (firejay5):

How would you graph it